OpenStudy (anonymous):

f(x) = [(x^3 + 9)]ln(x^3 + 9) whats the derivative?

OpenStudy (anonymous):

i think they want me to put it in this form: f' = A x^B + C x^D ln(x^F + G)

OpenStudy (anonymous):

since this is the product of two functions, let's use the product rule: [uv]' = u'v * v'u where u = x^3 + 9 v = ln (x^3 + 9) do you have u' and v'?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

I mean the derivative of u (du/dx) and the derivative of v (dv/dx)

OpenStudy (anonymous):

idk let me calculate it

OpenStudy (bahrom7893):

yup just use the product rule and then simplify.

OpenStudy (anonymous):

im having a problem so its deriv of x^3+9 times ln (x^3+9

OpenStudy (anonymous):

no,